Bed & Bath, a retailing company, has two departments—Hardware and Linens. The company’s most recent monthly contribution format income statement follows:

Department
Total Hardware Linens
Sales $ 4,230,000 $ 3,050,000 $ 1,180,000
Variable expenses 1,258,000 839,000 419,000
Contribution margin 2,972,000 2,211,000 761,000
Fixed expenses 2,250,000 1,440,000 810,000
Net operating income (loss) $ 722,000 $ 771,000 $ (49,000 )

A study indicates that $371,000 of the fixed expenses being charged to Linens are sunk costs or allocated costs that will continue even if the Linens Department is dropped. In addition, the elimination of the Linens Department will result in a 19% decrease in the sales of the Hardware Department.

Required:

What is the financial advantage (disadvantage) of discontinuing the Linens Department?Bed & Bath, a retailing company, has two departments-Hardware and Linens. The companys most recent monthly contribution form
Answer #1

Contribution margin for Hardware would be=2,211,000*(1-0.19)=1790910

Less:Fixed cost for Hardware=(1,440,000)

Net operating income for Hardware=350910

Less:Unavoidable fixed cost for Linens=(371,000)

Total net operating income would be=(20090)

Hence financial disadvantage=722,000-(20090)

=$742090